Pool problems:
A) Directions say to add 1 gallon of algae guard for every 50,000 gallons of water for the initial application. My pool currently contains 3,800 gallons. How much algae guard do I need to apply for the initial treatment of the pool in ounces?
B) Second part of instrutions say to add 1 quart per 10,000 gallons every 3-5 days. How much algae guard do I add to my pool every 3-5 days in ounces?

Answer by checkley75(3666)

SET UP A PAIR OF FRACTION SHOWING THE RELATIONSHIP OF BOTH MIXES THUS:
1/50000=X/3800 NOW WE CROSS MULTIPLY.
50000X=3800
X=3800/50000
X=.076 GALLONS ARE NEEDED. .076*8=.608 LBS. .608*16=9.728 OZ.
--------------------------------------------
1/10000=X/3800
10000X=3800
X=3800/10000
X=.38 QUARTS. .38*32=12.16 OUNCES.